# A measure-valued HJB perspective on Bayesian optimal adaptive control

### v1

@@ 3.1. Weak problem formulation
    (17) $\begin{split}\mathrm{d}\xi_{s}(\varphi)&=\Big{(}\xi_{s}\big{(}\varphi(\cdot)h(u_{s},\cdot)\big{)}-\xi_{s}\big{(}\varphi\big{)}\xi_{s}\big{(}h(u_{s},\cdot)\big{)}\Big{)}\mathrm{d}W_{s},\quad\varphi\in C_{b}.\end{split}$
@@ 3.2. Approximate problem formulation
[…]
@@
(19) $\mathrm{d}\xi_{t}(\varphi)$ $=\left(\xi_{t}\left(\varphi h(u_{t},\cdot)\right)-\xi_{t}\left(\varphi\right)\xi_{t}\left(h(u_{t},\cdot)\right)\right)\mathrm{d}W_{t},\;\varphi\in C_{b},\;\;\xi_{0}=\mu;$
since $\mu$ has compact support and $h(v,\cdot)$ is continuous for $v\in\mathcal{U}$ (see Lemma 4.1 below), existence and pathwise uniqueness of solutions to (19) is guaranteed by [36, Theorem 2.1]. For $\mu\in\mathcal{P}$ and $u\in\mathcal{A}^{n}$, $n\in\mathbb{N}$, we define
@@ 3.3. Associated HJB equation
$\sigma(v,\mu;\mathrm{d}x)=(h(v,x)-\mu(h(v,\cdot)))\mu(\mathrm{d}x).$
The HJB-equation associated with our problem is then given by (c.f. the dynamics in (17) and objective in (18)):
@@
for functions $u:\mathcal{P}\to\mathbb{R}$ we here make use of the notion of derivative as it was defined in [18, Section 4], which is essentially the same notion as the one referred to as the linear functional derivative in [12, Section 5.4].
@@ 3.4. Our main result
The functions $V^{+}$ and $V^{weak}$ effectively correspond to a piecewise constant approximation and a weak formulation of the separated problem associated with our main problem of study. Thanks to the above result, it is however possible to draw conclusions also about alternative formulations of the original problem itself: (i) Restricting in our main problem of study to controls which are piecewise constant (over arbitrary partitions) would not alter the value of the problem. (ii) Since the specifics of the underlying probability space do not have an impact on the value of the problem, considering a weak formulation where one also optimises over the underlying probability tuples, would still yield the same value. (iii) In the literature on stochastic control under partial information, it is common to consider so-called wide-sense controls which are adapted to a filtration which is larger than the observation filtration but still small enough for the filter to satisfy the associated filtering equation; allowing for controls of this type (adopting the definitions to the present context) would also not alter the value of the problem.
@@ 4. Stability properties of controlled MVMs
In this section we establish stability properties of our controlled MVMs with respect to the initial condition; these properties will become crucial in the upcoming analysis. We note that the proofs of these results do not make any use of the upcoming viscosity theory. Specifically, the proof of the first result relies on stability properties for classical SDEs, and the proof of the second result makes use of ideas similar to those used in the proof of [36, Theorem 2.1]; c.f. also the proof of [24, Theorem 3].
Before proceeding to the stability results, we establish a basic property of our action space:
@@
Let $\kappa<R$. Let $h(v,x)$ be given by (2) for $v\in\mathcal{U}$ and $x\in[-\kappa,\kappa]$. Then $h(v,\cdot)$ is continuous on $[-\kappa,\kappa]$ uniformly in $v\in\mathcal{U}$. In particular, $\{h(v,x):v\in\mathcal{U},x\in[-\kappa,\kappa]\}$ is a bounded subset of $\mathbb{R}$.
@@
For all $v\in\mathcal{U}$, $x\in[-\kappa,\kappa]$ and $n,m\in\mathbb{N}$,
@@
Since $\sum_{i=1}^{\infty}\left(\kappa/R\right)^{2i}$ converges, we obtain that $\sum_{i=1}^{\infty}v_{i}x^{i}$ is absolutely convergent, for all $x\in[-\kappa,\kappa]$ and $v\in\mathcal{U}$. Next, relying once again on (22), we have for every $\varepsilon>0$, that there exists some $n\in\mathbb{N}$, such that for all $x,y\in[-\kappa,\kappa]$ and $v\in\mathcal{U}$,
@@
[…]
@@
$(ii)$. Denote $\tilde{\mu}=\left(\mu(\{x_{1}\}),\ldots,\mu(\{x_{N}\})\right)$ and $\tilde{\nu}=\left(\nu(\{x_{1}\}),\ldots,\nu(\{x_{N}\})\right)$. Let $\tilde{\xi}^{u;\mu}$ and $\tilde{\xi}^{u;\nu}$ be solutions to (23) with initial conditions $\tilde{\mu}$ and $\tilde{\nu}$ respectively. Since the function $\tilde{\sigma}(v,\cdot)$ is uniformly Lipschitz for $v\in\mathcal{U}$, by [35, Corollary 2.5.5],
@@
Let $\mu\in\mathcal{P}$ and let $\mu_{k}$ be a sequence in $\mathcal{P}$ such that $\mathcal{W}(\mu_{k},\mu)\to 0$, as $k\to\infty$. Consider a filtered probability space supporting a Brownian motion $W$ and a sequence of progressively measurable $\mathcal{U}$-valued processes $u_{k}$, $k\in\mathbb{N}$, such that equation (17) admits pathwise unique solutions for $u=u_{k}$ and the initial conditions $\mu_{k}$ and $\mu$, respectively; denote the solutions by $\xi^{u_{k};\mu_{k}}$ and $\xi^{u_{k};\mu}$, $k\in\mathbb{N}$. Then, for any $t>0$,
@@
Let $u$ be a $\mathcal{U}$-valued progressively measurable process such that (17) admits pathwise unique solutions $\xi^{u;\mu_{k}}$ and $\xi^{u;\mu}$ with initial conditions $\mu_{k}$ and $\mu$, $k\in\mathbb{N}$. For ease of notation, we define
@@
and get a neater estimate
@@
Thus, we can do this iteratively to (26). If we regard (4) as the first iteration, then for the $m^{\text{th}}$ iteration, $m\geqslant 2$, $m\in\mathbb{N}$, we obtain the estimate
@@
Since $f\psi^{i}$ are Lipschitz for all $i\in\mathbb{N}$ and $\varepsilon$ is arbitrary, this means that every component in (30) converges to 00 as $k\to\infty$; since each $\Delta_{j}^{k}$ has finitely many components with finite coefficients, it follows that all the $\Delta_{j}^{k}$, $j=1,\ldots,m-1$, converge to 00 when $k\to\infty$. Therefore, from (29) we get
@@
[…]
@@
Note that $\{f_{i}\}_{i=1}^{\infty}$ strongly separates points in $\text{supp}(\mu_{0})$ and is closed under multiplication. Therefore, by [10, Theorem 6(a)], along this subsequence we have $\xi^{u_{k};\mu_{k}}\Rightarrow\xi^{u_{k};\mu}$ a.s. Since $\text{supp}(\mu_{0})$ is compact, this gives us the desired result. ∎
@@ 5. Viscosity theory and proof of main result
We first recall the definition of viscosity solutions which we make use of in this paper. To this end, let $C^{2}(\mathcal{P})$ denote the class of functions which are twice continuously differentiable in the sense of [18, Definition 4.7]. Further, recall that continuous MVMs have decreasing support in the sense that, with probability one, $\operatorname{supp}(\xi_{t})\subseteq\operatorname{supp}(\xi_{s})$, for $s\leq t$; see [18, Remark 2.3.(ii)]. Motivated by this fact, consider the partial order $\preceq$ defined on $\mathcal{P}$ by
@@ 5.1. DPP for piecewise constant controls
where we used that, by the standing Assumption 2.1, $\mu\mapsto k(\mu,v)$ is continuous, uniformly in $v\in\mathcal{U}$, and thus admits a concave and continuous modulus of continuity $w$ on the compact set $\mathcal{P}$, and the fact that $\mathcal{W}(\xi^{u_{k};\mu_{k}},\xi^{u_{k};\mu})$ is a submartingale for $k\in\mathbb{N}$ (c.f. [7, Proposition 4.1.(i)]).
@@
Next, we argue that a certain ‘pseudo Markov property’ holds within our context. To this end, let $n\in\mathbb{N}$ and recall the definition of $\delta_{n}$ from Section 3.2. Given a control $u\in\mathcal{A}^{n}$, we define a new control $u^{\delta_{n},\omega}\in\mathcal{A}^{n}$, for $\omega\in\Omega$, by
where
Lemma 5.3.
Let $n\in\mathbb{N}$ and $u\in\mathcal{A}^{n}$. Then, for every $\mu\in\mathcal{P}$ and $\mathbb{P}$-a.a. $\omega$,
$=\int_{\Omega}$ $\int_{0}^{\infty}e^{-\beta t}k\left(\xi^{u^{\delta_{n},\omega},\xi^{u;\mu}_{\delta_{n}}(\omega)}_{t}(\tilde{\omega}),u^{\delta_{n},\omega}_{t}(\tilde{\omega})\right)\mathrm{d}t\;\mathrm{d}\mathbb{P}(\tilde{\omega}).$
Thanks to the above continuity and pseudo Markov properties, we may establish the DPP following the arguments developed in [11]. In turn, by use of the DPP, we may prove existence of an $\varepsilon$-optimal strategy for the auxiliary problem. We present those results next; for completeness, we provide the full proofs in the present set-up in Appendix A.
@@ 5.2. Subsolution property of limiting problem
Making use of the fact that, for any $u\in C^{2}(\mathcal{P})$, the left-hand side of (20) is lower semicontinuous in $\mu$, and the fact that $V^{+}$ is upper semicontinuous on $\mathcal{P}$ (being the decreasing limit of continuous functions by Proposition 5.2), we note that $V^{+}$ is a subsolution of (20), if and only if, for any $\bar{\mu}\in\mathcal{P}$ and $\varphi\in C^{2}(\mathcal{P})$, one has the implication
[…]
Indeed, the ‘only if’ is immediate. Meanwhile, the ‘if’ follows by use of the same arguments as employed in the proof of [18, Lemma 6.6].
@@
Let $\bar{\mu}\in\mathcal{P}$ and $\varphi\in C^{2}(\mathcal{P})$ such that $\varphi(\bar{\mu})=V^{+}(\bar{\mu})$ and $V^{+}-\varphi<0$ on $D_{\bar{\mu}}\setminus\{\bar{\mu}\}$. Then, there exist sequences $m_{n}\to\infty$ and $\mu_{n}\to\bar{\mu}$, $\mu_{n}\in D_{\bar{\mu}}$, such that $\mu_{n}$ attains the maximum of $V^{m_{n}}-\varphi$ on $D_{\bar{\mu}}$ and $V^{m_{n}}(\mu_{n})\to V^{+}(\bar{\mu})$. In particular, $V^{m_{n}}\leq\varphi+\kappa_{n}$, where $\kappa_{n}:=(V^{m_{n}}-\varphi)(\mu_{n})\downarrow 0$.
@@
For $n\in\mathbb{N}$, let $N_{n}\geq n$ and $\varepsilon_{n}\leq 1/n$ such that, for all $m\geq N_{n}$, $\mu,\nu\in D_{\bar{\mu}}\cap B_{\varepsilon_{n}}(\bar{\mu})$,
@@
existence is guaranteed by Proposition 5.2. By the upper semicontinuity of $V^{+}$ (c.f. Proposition 5.2), there exists $\eta_{n}>0$ such that $V^{+}-\varphi\leq-2\eta_{n}$ on $D_{\bar{\mu}}\setminus B_{\varepsilon_{n}}(\bar{\mu})$. By possibly making $N_{n}$ even larger, we may ensure that, for all $m\geq N_{n}$, $\mu\in D_{\bar{\mu}}\setminus B_{\varepsilon_{n}}(\bar{\mu})$,
@@
Theorem 5.4 together with Lemma 5.6 immediately gives the following version of the DPP; we here focus on the inequality which will be needed for proving the subsolution property below and note that this inequality relies on the ‘deep inequality’ of the DPP:
@@
Let $\bar{\mu}\in\mathcal{P}$ and $\varphi\in C^{2}(\mathcal{P})$ such that $\varphi(\bar{\mu})=V^{+}(\bar{\mu})$ and $V^{+}-\varphi<0$ on $D_{\bar{\mu}}\setminus\{\bar{\mu}\}$. Then there exist sequences $m_{n}\to\infty$, $\mu_{n}\to\bar{\mu}$, $\mu_{n}\in D_{\bar{\mu}}$, and $\kappa_{n}\downarrow 0$, such that
$\varphi(\mu_{n})\leq\inf_{v\in\mathcal{U}}\mathbb{E}\left[\int_{0}^{\delta_{m_{n}}}e^{-\beta s}k\left(\xi^{v;\mu_{n}}_{s},v\right)\mathrm{d}s+e^{-\beta\delta_{m_{n}}}\varphi\left(\xi^{v;\mu_{n}}_{\delta_{m_{n}}}\right)\right]-\left(1-e^{-\beta\delta_{m_{n}}}\right)\kappa_{n}.$
@@
$\geq H\left(\bar{\mu},\varphi(\bar{\mu}),\dfrac{\partial^{2}\varphi}{\partial\mu^{2}}(\cdot,\cdot,\bar{\mu})\right).$
@@
Let $\mu\in\mathcal{P}$, $v\in\mathcal{U}$ and $\varphi\in C^{2}(\mathcal{P})$. By Lemma 4.1, $h(v,\cdot)$ is bounded on $\operatorname{supp}(\mu_{0})$. Hence, for any $f\in C_{b}$,
(36) $\int_{0}^{t}\left(\int_{\mathbb{R}}f(x)\left|h(v,x)-\xi^{v;\mu}_{s}(h(v,\cdot))\right|\xi^{v;\mu}_{s}(\mathrm{d}x)\right)^{2}\mathrm{d}s<\infty,\quad a.s.{}$
We may thus apply the Itô formula [18, Theorem 5.1] to obtain
@@
(37) $+\int_{0}^{t}e^{-\beta s}\int_{\mathbb{R}}\dfrac{\partial\varphi}{\partial\mu}(x;\xi^{v;\mu}_{s})\sigma_{s}(\mathrm{d}x)\mathrm{d}W_{s},$
where $\sigma(\mu,v;\mathrm{d}x)=(h(v,x)-\mu(h(v,\cdot)))\mu(\mathrm{d}x)$ and
$L\varphi(\mu,v)=\frac{1}{2}\int_{\mathbb{R}\times\mathbb{R}}\dfrac{\partial^{2}\varphi}{\partial\mu^{2}}(x,y;\mu)\sigma(\mu,v;\mathrm{d}x)\sigma(\mu,v;\mathrm{d}y).$
Making once again use of (36), we see that the last integral in (5.2) is in fact a martingale; we thus obtain
@@
Next, consider sequences $h_{k}\downarrow 0$ and $\mu_{k}\to\mu$; without loss of generality, suppose that $h_{k}\leq 1$, $k\in\mathbb{N}$. For $v\in\mathcal{U}$, by Lemma 4.1, $h(v,\cdot)$ is bounded and continuous on $\operatorname{supp}(\mu_{0})$ and therefore $L\varphi(\cdot,v)$ is continuous and thus admits a concave and continuous modulus of continuity, say $w$, on the compact set $\mathcal{P}$. Moreover, $\mathcal{W}(\xi^{v;\mu_{k}},\xi^{v;\mu})$ is a submartingale for $k\in\mathbb{N}$ (c.f. [7, Proposition 4.1.(i)]). Hence,

### v2

@@ 3.1. Weak problem formulation
    (18) $\begin{split}\mathrm{d}\xi_{s}(\varphi)&=\big(\xi_{s}\big(\varphi h(u_{s},\cdot)\big)-\xi_{s}(\varphi)\xi_{s}\big(h(u_{s},\cdot)\big)\big)\mathrm{d}W_{s},\quad\varphi\in C_{b}.\end{split}$
@@ 3.2. Approximate problem formulation
$\mathcal{A}^{n}=\Big\{u:[0,\infty)\times\Omega\to\mathcal{U}:u_{t}=\sum_{r\in\mathbb{T}^{n}}u_{r}\boldsymbol{1}_{[r,r+\delta_{n})}(t),\;\;t\in[0,\infty),\;\textrm{$u_{r}$ is $\mathcal{F}^{W}_{r}$-meas.}\Big\}.$
@@
(20) $\mathrm{d}\xi_{t}(\varphi)$ $=\big(\xi_{t}\big(\varphi h(u_{t},\cdot)\big)-\xi_{t}(\varphi)\xi_{t}\big(h(u_{t},\cdot)\big)\big)\mathrm{d}W_{t},\;\varphi\in C_{b},\;\;\xi_{0}=\mu;$
since $\mu$ has compact support and $h(v,\cdot)$ is continuous for $v\in\mathcal{U}$ (see Lemma 4.1 below), existence and pathwise uniqueness of solutions to (20) is guaranteed by [33, Theorem 2.1] (c.f. also [37, Theorems V.5 and V.6]). For $\mu\in\mathcal{P}$ and $u\in\mathcal{A}^{n}$, $n\in\mathbb{N}$, we define
@@ 3.3. Associated HJB equation
(21) $\sigma(v,\mu;\mathrm{d}x)=(h(v,x)-\mu(h(v,\cdot)))\mu(\mathrm{d}x).$
The HJB equation associated with our problem is then given by (c.f. the dynamics in (18) and objective in (19)):
@@
for functions $u:\mathcal{P}\to\mathbb{R}$ we here make use of the notion of derivative as it was defined in [16, Section 4], which is essentially the same notion as the one referred to as the linear functional derivative in [12, Section 5.4].
@@ 3.4. Our main result
The functions $V^{+}$ and $V^{weak}$ effectively correspond to a piecewise constant approximation and a weak formulation of the separated problem associated with our main problem of study. Thanks to the above result, it is however possible to draw conclusions also about alternative formulations of the original problem itself: (i) Restricting in our main problem of study to controls which are piecewise constant (over arbitrary partitions) would not alter the value of the problem. (ii) Since the specifics of the underlying probability space do not have an impact on the value of the problem, considering a weak formulation where one also optimises over the underlying probability tuples would still yield the same value. (iii) In the literature on stochastic control under partial information, it is common to consider so-called wide-sense controls which are adapted to a filtration which is larger than the observation filtration but still small enough for the filter to satisfy the associated filtering equation; allowing for controls of this type (adopting the definitions to the present context) would also not alter the value of the problem.
@@ 4. Stability properties of controlled MVMs
In this section we prove stability properties of our controlled MVMs with respect to the initial condition; these properties will become crucial in the upcoming analysis. Before proceeding to the key results, we verify some basic properties of our action space:
@@
The set $\{h(v,x):v\in\mathcal{U},x\in\operatorname{supp}(\mu_{0})\}$ is a bounded subset of $\mathbb{R}$ and the function $h(v,\cdot)$ is continuous on $\operatorname{supp}(\mu_{0})$ uniformly in $v\in\mathcal{U}$.
@@
Let $0<\kappa<R$ such that $\operatorname{supp}(\mu_{0})\subseteq[-\kappa,\kappa]$. Further, let $v\in\mathcal{U}$ and $x\in\operatorname{supp}(\mu_{0})$. Then, for all $n,m\in\mathbb{N}$,
@@
Since $\sum_{i=0}^{\infty}\left(\kappa/R\right)^{2i}$ converges, we obtain that $\sum_{i=0}^{\infty}v_{i}x^{i}$ is absolutely convergent. A similar calculation yields the boundedness claim. Next, relying once again on (24), we have that for every $\varepsilon>0$, there exists some $n\in\mathbb{N}$, such that for all $v\in\mathcal{U}$ and $x,y\in\operatorname{supp}(\mu_{0})$,
@@
Our first stability result concerns finitely supported MVMs; its proof relies on stability properties for classical SDEs.
@@
    where $\xi^{u;\mu}$ and $\xi^{u;\nu}$ denote the respective solutions to equation (18) and $\|\cdot\|$ is the Euclidean norm.
@@
$(ii)$. Denote $\tilde{\mu}=\left(\mu(\{x_{1}\}),\ldots,\mu(\{x_{N}\})\right)$ and $\tilde{\nu}=\left(\nu(\{x_{1}\}),\ldots,\nu(\{x_{N}\})\right)$. Let $\tilde{\xi}^{u;\mu}$ and $\tilde{\xi}^{u;\nu}$ be solutions to (25) with initial conditions $\tilde{\mu}$ and $\tilde{\nu}$ respectively. Since the function $\tilde{\sigma}(v,\cdot)$ is uniformly Lipschitz for $v\in\mathcal{U}$, by [32, Corollary 2.5.5],
@@
Our second stability result concerns MVMs with general initial conditions. Its proof makes use of ideas similar to those employed in the proof of [33, Theorem 2.1] (c.f. also the proof of [23, Theorem 3]). The convergence is notably uniform in the control; this is needed for the upcoming proofs of Proposition 5.2 and the continuity of $V^{+}$ (c.f. (43)).
@@
Let $\mu\in\mathcal{P}$ and let $\mu_{k}$ be a sequence in $\mathcal{P}$ such that $\mathcal{W}(\mu_{k},\mu)\to 0$ as $k\to\infty$. Consider a filtered probability space supporting a Brownian motion $W$ and a sequence of progressively measurable $\mathcal{U}$-valued processes $u_{k}$, $k\in\mathbb{N}$. Suppose that $\xi^{u_{k};\mu_{k}}$ and $\xi^{u_{k};\mu}$, $k\in\mathbb{N}$, are solutions to equation (18) for $u=u_{k}$ and the initial conditions $\mu_{k}$ and $\mu$, respectively. Then, the solutions are unique (up to indistinguishability) and for any $t>0$,
@@
Let $u$ be a $\mathcal{U}$-valued progressively measurable process such that (18) admits solutions $\xi^{u;\mu_{k}}$ and $\xi^{u;\mu}$ with initial conditions $\mu_{k}$ and $\mu$, $k\in\mathbb{N}$. For ease of notation, we define
@@
Since (18) is indifferent with respect to parallel shifts in $h(v,\cdot)$, without loss of generality, we may assume that $v_{0}\equiv 0$ for all $v\in\mathcal{U}$.
@@
which allows us to rewrite the above estimate as
@@
We do this iteratively to (28). If we regard (4) as the first iteration, then for the $m^{\text{th}}$ iteration, $\geqslant 2, mm\in\mathbb{N}$, we obtain the estimate
@@
Since $f\psi^{i}$ are Lipschitz for all $i\in\mathbb{N}$ and $\varepsilon$ is arbitrary, this means that every component in (32) converges to 0 as $k\to\infty$; since each $\Delta_{j}^{k}$ has finitely many components with finite coefficients, it follows that all the $\Delta_{j}^{k}$, $j=1,\ldots,m-1$, converge to 0 when $k\to\infty$. Therefore, from (31) we get
@@
$\left|\xi^{u_{k};\mu_{k}}_{t}(f)-\xi^{u_{k};\mu}_{t}(f)\right|\xrightarrow[k\to\infty]{}0\quad\text{a.s.}$
@@
Note that $\{f_{i}\}_{i=1}^{\infty}$ is an algebra which strongly separates points in $\text{supp}(\mu_{0})$. Therefore, by [22, Theorem 3.4.5(b)], along this subsequence we have $\xi^{u_{k};\mu_{k}}\Rightarrow\xi^{u_{k};\mu}$ a.s. Since $\text{supp}(\mu_{0})$ is compact, this gives us the desired result.
Step 3: We show pathwise uniqueness of solutions to (18). Let $\xi^{u;\mu}$ and $\tilde{\xi}^{u;\mu}$ be two solutions to (18) defined on the same probability space and with respect to the same Brownian motion $W$, control $u$ and initial condition $\mu$. Define, for $f\in C_{b}$,
$\rho_{t}(f)=\mathbb{E}\left|\xi^{u;\mu}_{t}(f)-\tilde{\xi}^{u;\mu}_{t}(f)\right|^{2}.$
Using similar arguments and notation as in step 1, we have
$\rho_{t}(f)\leq M\int_{0}^{t}\left(\sum_{i=1}^{\infty}\rho_{s}(f\psi^{i})+\|f\|^{2}\sum_{i=1}^{\infty}\rho_{s}(\psi^{i})+\rho_{s}(f)\right)\mathrm{d}s.$
By applying the iteration procedure (note that it simplifies considerably since all $D_{k}$ vanish), we obtain that for any $m\in\mathbb{N}$,
$\rho_{t}(f)\leq 4M^{m}(2L+1)^{m}\frac{t^{m}}{m!}\|f\|^{2}.$
Sending $m\to\infty$, we get $\rho_{t}(f)=0$ for all $t\geq 0$ and $f\in C_{b}$. By arguments similar to those used in step 2, using the countable sequence $\{f_{i}\}_{i=1}^{\infty}$ and [22, Theorem 3.4.5(a)], we get $\xi^{u;\mu}_{t}=\tilde{\xi}^{u;\mu}_{t}$ a.s. for all $t\geq 0$. Moreover, by [16, Remark 2.3], the trajectories of $\xi^{u;\mu}$ and $\tilde{\xi}^{u;\mu}$ are continuous. Hence, the two processes are indistinguishable, completing the proof. ∎
@@ 5. Viscosity theory and proof of main result
We first recall the definition of viscosity solutions which we make use of in this paper. To this end, let $C^{2}(\mathcal{P})$ denote the class of functions which are twice continuously differentiable in the sense of [16, Definition 4.7]. Further, recall that continuous MVMs have decreasing support in the sense that, with probability one, $\operatorname{supp}(\xi_{t})\subseteq\operatorname{supp}(\xi_{s})$, for $s\leq t$; see [16, Remark 2.3.(ii)]. Motivated by this fact, consider the partial order $\preceq$ defined on $\mathcal{P}$ by
@@ 5.1. DPP for piecewise constant controls
where we used that, by the standing Assumption 2.1, $\mu\mapsto k(\mu,v)$ is continuous, uniformly in $v\in\mathcal{U}$, and thus admits a concave and continuous modulus of continuity $w$ on the compact set $\mathcal{P}$, and the fact that $\mathcal{W}(\xi^{u_{k};\mu_{k}},\xi^{u_{k};\mu})$ is a submartingale for $k\in\mathbb{N}$ (c.f. [8, Proposition 4.1.(i)]).
@@
Thanks to the above continuity properties, we may establish the DPP following the arguments developed in [11]. By use of the DPP, we may then prove existence of an $\varepsilon$-optimal strategy for the auxiliary problem. We present those results next; for completeness, the proofs are provided in Appendix A.
@@ 5.2. Subsolution property of limiting problem
Making use of the fact that, for any $u\in C^{2}(\mathcal{P})$, the left-hand side of (22) is lower semicontinuous in $\mu$, and the fact that $V^{+}$ is upper semicontinuous on $\mathcal{P}$ (being the decreasing limit of continuous functions by Proposition 5.2), we also note that $V^{+}$ is a subsolution of (22), if and only if, for any $\bar{\mu}\in\mathcal{P}$ and $\varphi\in C^{2}(\mathcal{P})$, one has the implication
(36) $\text{$\varphi(\bar{\mu})=V^{+}(\bar{\mu})$ and $\varphi>V^{+}$ on $D_{\bar{\mu}}\setminus\{\bar{\mu}\}$}\;\;\Longrightarrow\;\;H\left(\bar{\mu},\varphi(\bar{\mu}),\dfrac{\partial^{2}\varphi}{\partial\mu^{2}}(\cdot,\cdot;\bar{\mu})\right)\leq 0.$
Indeed, the ‘only if’ is immediate. Meanwhile, the ‘if’ follows by use of the same arguments as employed in the proof of [16, Lemma 6.6].
@@
Let $\bar{\mu}\in\mathcal{P}$ and $\varphi\in C^{2}(\mathcal{P})$ such that $\varphi(\bar{\mu})=V^{+}(\bar{\mu})$ and $V^{+}-\varphi<0$ on $D_{\bar{\mu}}\setminus\{\bar{\mu}\}$. Then, there exist sequences $m_{n}\to\infty$ and $\mu_{n}\to\bar{\mu}$, $\mu_{n}\in D_{\bar{\mu}}$, such that $\mu_{n}$ attains the maximum of $V^{m_{n}}-\varphi$ on $D_{\bar{\mu}}$ and $V^{m_{n}}(\mu_{n})\to V^{+}(\bar{\mu})$.
In particular, defining $\kappa_{n}:=(V^{m_{n}}-\varphi)(\mu_{n})\geq 0$ then $\kappa_{n}\to 0$ and $V^{m_{n}}\leq\varphi+\kappa_{n}$ on $D_{\bar{\mu}}$.
@@
We first note that $D_{\bar{\mu}}$ is a closed subset of the compact set $\mathcal{P}$; hence, $D_{\bar{\mu}}$ is itself a compact set. For $n\in\mathbb{N}$, let $N_{n}\geq n$ and $\varepsilon_{n}\leq 1/n$ such that, for all $m\geq N_{n}$, $\mu,\nu\in D_{\bar{\mu}}\cap B_{\varepsilon_{n}}(\bar{\mu})$,
@@
existence is guaranteed by the fact that $V^{n}$, $n\in\mathbb{N}$, is continuous and decreasing in $n$ (c.f. Proposition 5.2). By the upper semicontinuity of $V^{+}$ (c.f. again Proposition 5.2), there exists $\eta_{n}>0$ such that $V^{+}-\varphi\leq-2\eta_{n}$ on $D_{\bar{\mu}}\setminus B_{\varepsilon_{n}}(\bar{\mu})$. By possibly making $N_{n}$ even larger, we may ensure that, for all $m\geq N_{n}$, $\mu\in D_{\bar{\mu}}\setminus B_{\varepsilon_{n}}(\bar{\mu})$,
@@
Theorem 5.3 together with Lemma 5.5 immediately gives the following version of the DPP (note that the full strength of the lemma is in fact not used); we here focus on the inequality which will be needed for proving the subsolution property below and note that this inequality relies on the ‘deep inequality’ of the DPP:
@@
Let $\bar{\mu}\in\mathcal{P}$ and $\varphi\in C^{2}(\mathcal{P})$ such that $\varphi(\bar{\mu})=V^{+}(\bar{\mu})$ and $V^{+}-\varphi<0$ on $D_{\bar{\mu}}\setminus\{\bar{\mu}\}$. Then there exist sequences $m_{n}\to\infty$, $\mu_{n}\to\bar{\mu}$, $\mu_{n}\in D_{\bar{\mu}}$, such that
$\varphi(\mu_{n})\leq\inf_{v\in\mathcal{U}}\mathbb{E}\left[\int_{0}^{\delta_{m_{n}}}e^{-\beta s}k\left(\xi^{v;\mu_{n}}_{s},v\right)\mathrm{d}s+e^{-\beta\delta_{m_{n}}}\varphi\left(\xi^{v;\mu_{n}}_{\delta_{m_{n}}}\right)\right].$
@@
$\geq H\left(\bar{\mu},\varphi(\bar{\mu}),\dfrac{\partial^{2}\varphi}{\partial\mu^{2}}(\cdot,\cdot;\bar{\mu})\right).$
@@
Let $\mu\in\mathcal{P}$, $v\in\mathcal{U}$ and $\varphi\in C^{2}(\mathcal{P})$. By Lemma 4.1, $h(v,\cdot)$ is bounded on $\operatorname{supp}(\mu_{0})$. Hence, for any $f\in C_{b}$, we have $\mathbb{P}\otimes\mathrm{d}t$-a.e.,
(38) $\int_{0}^{t}\left(\int_{\mathbb{R}}f(x)\left|h(v,x)-\xi^{v;\mu}_{s}(h(v,\cdot))\right|\xi^{v;\mu}_{s}(\mathrm{d}x)\right)^{2}\mathrm{d}s<\infty.$
We may thus apply the Itô formula [16, Theorem 5.1] to obtain, for $t\geq 0$,
@@
(39) $+\int_{0}^{t}e^{-\beta s}\int_{\mathbb{R}}\dfrac{\partial\varphi}{\partial\mu}(x;\xi^{v;\mu}_{s})\sigma(v,\xi^{v;\mu}_{s};\mathrm{d}x)\mathrm{d}W_{s},$
where $\sigma(v,\mu;\mathrm{d}x)$ is given by (21) and
$L\varphi(\mu,v)=\frac{1}{2}\int_{\mathbb{R}\times\mathbb{R}}\dfrac{\partial^{2}\varphi}{\partial\mu^{2}}(x,y;\mu)\sigma(v,\mu;\mathrm{d}x)\sigma(v,\mu;\mathrm{d}y).$
Making again use of (38), we see that the last integral in (5.2) is in fact a martingale; we thus obtain
@@
Next, consider sequences $h_{k}\downarrow 0$ and $\mu_{k}\to\mu$; without loss of generality, suppose that $h_{k}\leq 1$, $k\in\mathbb{N}$. For $v\in\mathcal{U}$, by Lemma 4.1, $h(v,\cdot)$ is continuous on $\operatorname{supp}(\mu_{0})$ and therefore $L\varphi(\cdot,v)$ is continuous and thus admits a concave and continuous modulus of continuity, say $w$, on the compact set $\mathcal{P}$. Moreover, $\mathcal{W}(\xi^{v;\mu_{k}},\xi^{v;\mu})$ is a submartingale for $k\in\mathbb{N}$ (c.f. [8, Proposition 4.1.(i)]). Hence,